# A New Causal Rule Learning Approach to Interpretable Estimation of Heterogeneous Treatment Effect

### v2

@@ 4 Evaluation of CRL Performance
In this section, we use both simulated data and real-world data to evaluate how well CRL performs in estimating and providing insights to HTE. Specifically, the simulated RCT data is mainly performed to evaluate the estimation accuracy of CRL while the real-world application demonstrates how to apply CRL in observational data and how CRL gives insights to us on real-world problems (enhanced understanding). In addition, we also compare CRL with several previously mentioned baseline methods, namely outcome weighted estimation (OWE) [50], causal tree (CT) [3], causal forest (CF) [45], patient rule induction method (PRIM) [9] and causal rule ensemble (CRE) [28] in both the simulation experiments and real-world applications. Baseline methods are implemented with R package personalized [24] for OWE (linear), causalTree [4] for causal tree, subgrpID [22] for PRIM, grf [40] for causal forest.
4.1 Simulation study
@@ 4.1.1 Data setups
To evaluate how well CRL performs in estimating HTE, we first generate simulated RCT data sets $\{X_{i},A_{i},Y_{i}\}_{i=1}^{N}$ based on assumption (2) with varying configurations. Specifically, we generate covariates $X\in\mathbb{R}^{9}$ where the first three elements (features) $x_{1},x_{2}$, and $x_{3}$ are randomly drawn from $\{0,1\}$ with equal probability ($x_{3}$ is sampled differently for some setups of the true group, shown later). The last six features are $\{x_{j}|j=4,5,\dots,9\}\overset{\mathrm{i.i.d}}{\sim}\mathcal{N}(0,2)$. For simplicity, we ignore the sample subscript notation $i$, and the remaining subscript represents certain features unless otherwise noted. The baseline effect is defined as $f(X)=1+x_{1}^{2}+3x_{1}x_{6}+0.4x_{4}x_{7}+x_{5}-x_{6}+0.5x_{3}x_{8}$. Based on (2), we can define any number of true subgroups and their related true CATEs via a rule form (shown later). Once we set the true $\tau(X)$, we define the potential outcomes for both the negative and positive treatments: $Y(-1)=f(X)+\epsilon$ and $Y(1)=Y(-1)+\tau(X)$, where $\epsilon\sim\mathcal{N}(0,1)$. We imitate the simplest RCT assignment of the treatment arm: $A_{i}$ is sampled from $\{-1,1\}$ with equal probability. Therefore, the observed outcome in our data set is $Y_{i}=2^{-1}(1+A_{i})Y_{i}(1)+2^{-1}(1-A_{i})Y_{i}(-1)$.
To sum up, the observed data set $\{X_{i},A_{i},Y_{i}\}^{N}_{i=1}$, together with the predefined baseline effect $\{f(X_{i})\}^{N}_{i=1}$ and true ITE $\{\tau(X_{i})\}^{N}_{i=1}$ forms the basis of the ground truth of the data generation. To avoid randomness and to test how the performance of CRL changes with the data settings, we make a few variations to our simulated data:
@@
Together, we have $3\times 3\times 6=54$ data sets. We first make 100 re-sampled repetitions for each data set and randomly split each repetition into a 70% training set and 30% test set. Then, we apply the overall CRL workflow as well as other baseline methods on the training data and report performance on the test data, averaging over 100 repetitions.
@@ 4.1.2 Performance metrics for method evaluation
Performance metrics used in this study focus on how accurately a method estimates ITE. To evaluate so, we use Mean Squared Error (MSE), i.e., $MSE(\hat{\tau})=N^{-1}\sum^{N}_{i=1}(\hat{\tau}(X_{i})-\tau(X_{i}))^{2}$, where $\hat{\tau}(X_{i})$ and $\tau(X_{i})$ are a certain estimate and the true ITE of instance $i$ respectively. The smaller the value, the better an estimator does. Besides MSE, we also evaluate performance from two other supplementary perspectives, namely Mean Potential Outcomes (MPO) and Population Overlap (PO).
MPO, defined as $N^{-1}\sum^{N}_{i=1}{Y_{i}(d(X_{i}))}$, measures how good the average outcomes can be under a certain treatment assignment strategy $d(X)$. The larger the value of MPO, the better the treatment strategy is. For an intuitive comparison, we use the simplest strategy that the treatment recommendation is purely based on the sign of ITE estimates, i.e., $d(X_{i})=1$ if $\hat{\tau}(X_{i})>0$ and $-1$ otherwise.
We define $PO={|T\cap I|}/{|T\cup I|}\in[0,1]$ where $T=\{i\in N|\tau(X_{i})>0\}$ and $I=\{i\in N|\hat{\tau}(X_{i})>0\}$ represent respectively the sets of instances whose true ITE and estimated ITE are positive and ||$\cdot$|| ⋅ | is the number of instances in a set. This measure evaluates the degree of intersection of the two sets. When $T\cap I=\emptyset$, $PO=0$ and when $T=I$, $PO=1$. Larger values are preferred.
We also use the above metrics for tuning the parameters involved in CRL and other methods. Please see Appendix C for details.
@@ 4.1.3 Simulation results
Figure 4: Performance comparison of CRL (black) and baseline methods CT (blue), CF (green), OWE (yellow) applied on the simulated data. All performance metrics are averaged over 100 repetitions for each data set. (a) Mean squared error of treatment effect. (b) Mean potential outcomes. Note that CRL, CT and CF have identical performance on MPO, hence their curves overlap. (c) Mean population overlap. (d) Proportion of fake CF rules filtered out by CRL.
@@
Figure 5: Comparison of mean squared error of treatment effect on specific subpopulations between CRL and baseline methods CRE (pink) and PRIM (red) applied on the simulated data. Each value shown is the average over 100 repetitions for each data set. (a) MSE of CRL versus CRE on the subgroup identified by CRE. (b) MSE of CRL versus PRIM on the subgroup identified by PRIM.
@@
This section shows the results of the performance comparison between CRL and other baseline methods for each of the 54 data sets (setting). Since the baseline methods vary in research focus and model outputs, not all methods can be directly compared. Specifically, CT, CF, OWE and CRL are able to give ITE estimates for every instance in the data, so we compare and show their estimation accuracy on the whole population in Figure 4. For CRE and PRIM, which only identify specific subpopulations with enhanced treatment effect, we compare them with CRL only on the subgroups they respectively identify, shown in Figure 5
@@
Figure 4(a) shows the comparison of MSE between CT (blue), CF (green), OWE (yellow), and CRL (black). For panels from the top to bottom, the number of observations of the data set varies from 2,500 to 10,000, and from left to right, the number of true positive subgroups varies from one to five. Effect size ($k$) is reflected in the x-axis and each value shown is an average of 100 runs on the specific setting. Figure 4(b), 4(c), 4(d), 5(a) and 5(b) are presented in the same way as Figure 4(a) for other metrics.
@@
Figure 4(a) shows that each method’s MSE increases as the effect size grows across different configurations. In general, CRL is almost always among the best performers across all settings. For cases where $num.true.grp=1$, CT, CRL, and OWE have almost the same performance. Hence their curves overlap. Obviously, the superiority of CRL increases gradually as the number of observations and true subgroups increase under large effect sizes (e.g., $k\in\{20,25,30\}$). Besides, from Figure 5, we see that no matter for the subgroups identified by CRE (pink) or PRIM (red), CRL has consistently much lower MSE, indicating that CRL performs better for treatment effect estimation of complex diseases which have many potential contributing factors and interactions between these factors. In contrast, both CRE and PRIM identify a limited number of subgroups (most of the time they only identify one single subgroup) and variables, failing to achieve the goal.
@@
Figure 4(b) compares how well the involved methods do in treatment recommendations using MPO. CRL, CT, and CF show almost the same good performance on MPO and outperform OWE in all settings. The identical performance of the former three methods is a result of our treatment assignment strategy: We recommend treatment based solely on the sign of ITE estimates. Regardless of how these methods differ in the magnitude of the estimates, they yield identical treatment recommendations as long as their estimates have identical signs, hence the same values on MPO. Figure 4(c) shows that CRL has comparable performance to baseline methods with regard to PO across most settings (with higher overlap value and lower variance), indicating that it is able to recover the majority of populations with positive treatment effect.
@@
Since CRL filters out rules from the original rule set generated by CF, we also compute the proportion of CF rules filtered out by CRL, i.e., $1-{W_{CRL}}/{W_{CF}}$ where $W_{CRL}=|\{r_{m}|r_{m}\in I_{CRL},r_{m}\notin T\}|$ represents the number of fake rules in the rule set ($I_{CRL}$) identified by CRL, and $W_{CF}$ represents the same for CF. The bigger the value, the better CRL filters out fake rules. As indicated in Figure 4(d), through the additional D-learning selection and rule analysis procedure, CRL is able to filter out a high proportion of fake rules CF generates.

### v3

@@ 4 Evaluation of CRL Performance
In this section, we use both simulated data and real-world data to evaluate how well CRL performs in estimating and providing insights to HTE. Specifically, the simulated RCT data is mainly used to evaluate the estimation accuracy of CRL, while the real-world application demonstrates how to apply CRL in observational data and how CRL gives insights to us on real-world problems (enhanced understanding). In addition, we also compare CRL with several previously mentioned baseline methods, namely outcome weighted estimation (OWE) [zhao2012estimating], causal tree (CT) [athey2016recursive], causal forest (CF) [wager2018estimation], patient rule induction method (PRIM) [chen2015prim] and causal rule ensemble (CRE) [bargaglistoffi2024causal] in both the simulation experiments and real-world applications. Baseline methods are implemented with R package personalized [R_personalized] for OWE (linear), causalTree [causalTree_pkg] for causal tree, subgrpID [subgrp2017huang] for PRIM, grf [tibshirani2023package] for causal forest. Code is available at https://github.com/yingwu2017/Causal-Rule-Learning/.
4.1 Simulation study one: the overall performance of CRL
@@ 4.1.1 Data setups
To evaluate how well CRL performs in estimating HTE, we first generate simulated RCT data sets $\{X_{i},A_{i},Y_{i}\}_{i=1}^{N}$ based on assumption (2) with varying configurations. Specifically, we generate covariates $X\in\mathbb{R}^{9}$ where the first three elements (features) $x_{1},x_{2}$, and $x_{3}$ are randomly drawn from $\{0,1\}$ with equal probability ($x_{3}$ is sampled differently for some setups of the true group, shown later). The last six features are $\{x_{j}|j=4,5,\dots,9\}\overset{\mathrm{i.i.d}}{\sim}\mathcal{N}(0,2)$. For simplicity, we ignore the sample subscript notation $i$, and the remaining subscript represents certain features unless otherwise noted. The baseline effect is defined as $f(X)=1+x_{1}^{2}+3x_{1}x_{6}+0.4x_{4}x_{7}+x_{5}-x_{6}+0.5x_{3}x_{8}$. Based on (2), we can define any number of true subgroups and their related true CATEs via a rule form (shown later). Once we set the true $\tau(X)$, we define the potential outcomes for both the negative and positive treatments: $Y(-1)=f(X)+\epsilon$ and $Y(1)=Y(-1)+\tau(X)$, where $\epsilon\sim\mathcal{N}(0,1)$. We simulate the simplest RCT assignment of the treatment arm: $A_{i}$ is sampled from $\{-1,1\}$ with equal probability. Therefore, the observed outcome in our data set is $Y_{i}=2^{-1}(1+A_{i})Y_{i}(1)+2^{-1}(1-A_{i})Y_{i}(-1)$.
To sum up, the observed data set $\{X_{i},A_{i},Y_{i}\}^{N}_{i=1}$, together with the predefined baseline effect $\{f(X_{i})\}^{N}_{i=1}$ and true ITE $\{\tau(X_{i})\}^{N}_{i=1}$ forms the basis of the ground truth of the data generation. To avoid randomness and to test how the performance of CRL changes with the data settings, we introduce a few variations to our simulated data:
@@
Together, we have $3\times 3\times 6=54$ data sets. We first make 100 resampled repetitions for each data set and randomly split each repetition into a 70% training set and a 30% test set. Then, we apply the overall CRL workflow as well as other baseline methods on the training data and report performance on the test data, averaging over 100 repetitions.
@@ 4.1.2 Performance metrics for method evaluation
Performance metrics used in this study focus on how accurately a method estimates ITE. To evaluate so, we use Mean Squared Error (MSE), i.e., $MSE(\hat{\tau})=N^{-1}\sum^{N}_{i=1}(\hat{\tau}(X_{i})-\tau(X_{i}))^{2}$, where $\hat{\tau}(X_{i})$ and $\tau(X_{i})$ are a certain estimate and the true ITE of instance $i$ , respectively. The smaller the value, the better an estimator does. Besides MSE, we also evaluate performance from two other supplementary perspectives, namely Mean Potential Outcomes (MPO) and Population Overlap (PO).
MPO, defined as $N^{-1}\sum^{N}_{i=1}{Y_{i}(d(X_{i}))}$, measures how good the average outcomes can be under a certain treatment assignment strategy $d(X)$. The larger the value of MPO, the better the treatment strategy is. For an intuitive comparison, we use the simplest strategy that the treatment recommendation is purely based on the sign of ITE estimates, i.e., $d(X_{i})=1$ if $\hat{\tau}(X_{i})>0$ and 1-1 otherwise.
We define $PO={|T\cap I|}/{|T\cup I|}\in[0,1]$ where $T=\{i\in N|\tau(X_{i})>0\}$ and $I=\{i\in N|\hat{\tau}(X_{i})>0\}$ represent respectively the sets of instances whose true ITE and estimated ITE are positive and $|\cdot|$ is the number of instances in a set. This measure evaluates the degree of intersection of the two sets. When $T\cap I=\emptyset$, $PO=0$ and when $T=I$, $PO=1$. Larger values are preferred.
We also use the above metrics for tuning the parameters involved in CRL and other methods. Please see Appendix D for details.
@@ 4.1.3 Simulation results
Figure 4: Performance comparison of CRL (black asterisk) and baseline methods CT (blue circle), CF (green square), OWE (yellow triangle) applied on the simulated data. All performance metrics are averaged over 100 repetitions for each data set. (a) Mean-squared error of treatment effect. (b) Mean potential outcomes. Note that CRL, CT, and CF have identical performance on MPO, hence their curves overlap. (c) Mean population overlap. (d) Proportion of fake CF rules filtered out by CRL.
@@
Figure 5: Comparison of mean squared error of treatment effect on specific subpopulations between CRL and baseline methods CRE (pink cross) and PRIM (red X) applied on the simulated data. Each value shown is the average over 100 repetitions for each data set. (a) MSE of CRL versus CRE on the subgroup identified by CRE. (b) MSE of CRL versus PRIM on the subgroup identified by PRIM.
@@
This section shows the results of the performance comparison between CRL and other baseline methods for each of the 54 data sets (settings). Since the baseline methods vary in research focus and model outputs, not all methods can be directly compared. Specifically, CT, CF, OWE, and CRL are able to give ITE estimates for every instance in the data, so we compare and show their estimation accuracy on the whole population in Figure 4. For CRE and PRIM, which only identify specific subpopulations with enhanced treatment effect, we compare them with CRL only on the subgroups they respectively identify, shown in Figure 5
@@
Figure 4(a) shows the comparison of MSE between CT (blue), CF (green), OWE (yellow), and CRL (black). For panels from the top to bottom, the number of observations of the data set varies from 2,500 to 10,000, and from left to right, the number of true positive subgroups varies from one to five. Effect size ($k$) is reflected on the x-axis, and each value shown is an average of 100 runs on the specific setting. Figure 4(b), 4(c), 4(d), 5(a) and 5(b) are presented in the same way as Figure 4(a) for other metrics.
@@
Figure 4(a) shows that each method’s MSE increases as the effect size grows across different configurations. In general, CRL is almost always among the best performers across all settings. For cases where $num.true.grp=1$, CT, CRL, and OWE have almost the same performance. Hence, their curves overlap. Obviously, the superiority of CRL increases gradually as the number of observations and true subgroups increase under large effect sizes (e.g., $k\in\{20,25,30\}$). Besides, from Figure 5, we see that no matter for the subgroups identified by CRE (pink) or PRIM (red), CRL has a consistently much lower MSE, indicating that CRL performs better for treatment effect estimation of complex diseases that have many potential contributing factors and interactions between these factors. In contrast, both CRE and PRIM identify a limited number of subgroups (most of the time, they only identify a single subgroup) and variables, failing to achieve the goal.
@@
Figure 4(b) compares how well the involved methods do in treatment recommendations using MPO. CRL, CT, and CF show almost the same good performance on MPO and outperform OWE in all settings. The identical performance of the former three methods is a result of our treatment assignment strategy: We recommend treatment solely on the sign of ITE estimates. Regardless of how these methods differ in the magnitude of the estimates, they yield identical treatment recommendations as long as their estimates have identical signs, hence the same values on MPO. However, in practice where decisions incorporate cost-benefit analysis or patient prioritization based on the expected magnitude of benefit (as reflected by the ITE magnitude), the superior accuracy of CRL’s ITE estimates (Figure 4(a)) would be critical for developing more nuanced and effective strategies.
Figure 4(c) shows that CRL has comparable performance to baseline methods concerning PO across most settings (with higher overlap value and lower variance), indicating that it can recover the majority of populations with positive treatment effects.
@@
Since CRL filters out rules from the original rule set generated by CF, we also compute the proportion of CF rules filtered out by CRL, i.e., $1-{W_{CRL}}/{W_{CF}}$ where $W_{CRL}=|\{r_{m}|r_{m}\in I_{CRL},r_{m}\notin T\}|$ represents the number of fake rules in the rule set ($I_{CRL}$) identified by CRL, and $W_{CF}$ represents the same for CF. The bigger the value, the better CRL filters out fake rules. As indicated in Figure 4(d), through the additional D-learning selection and rule analysis procedure, CRL can filter out a high proportion of fake rules CF generates.
At the end of this section, we remind the reader that all variables in the simulated data used herein are independently sampled. To evaluate the performance of CRL under more realistic conditions where variables exhibit correlations, we generate multiple datasets with varying degrees of correlation to test the robustness of CRL’s performance in the overall analysis and the p-values and their interpretation in the rule decomposition analysis under correlated settings. Results from both experiments confirm that CRL exhibits strong robustness in both performance metrics and rule interpretation, even under significant variable correlations. For detailed results and discussions, please refer to Appendix E.
4.2 Simulation study two: Effectiveness of D-learning in rule selection
In CRL’s workflow, the D-learning method plays a core role by selecting a sparse and refined rule set from a candidate rule set that includes the true rules and fake rules with corresponding subgroup CATEs, which greatly enhances the generalization and interpretability of the proposed method. Therefore, this study mainly tests D-learning from two aspects: i) The adaptability of D-learning in the rule selection step. Precisely, is D-learning able to adaptively select an appropriate number of rules according to the underlying treatment heterogeneity and data complexity? We expect the trend that when the number of true subgroups ($M^{*}$) increases with the number of selected (non-zero) rules. ii) The accuracy of D-learning in the discovery of true rules, i.e., how many true rules can be recovered via D-learning? Higher values are preferred.
4.2.1 Data setups
Instead of generating explicit covariates and designing specific ground-truth rules as we do in study one, this study starts with directly generating the data fed to the D-learning method, i.e., $\{r_{m},\tau_{m}\}^{M}_{m=1}$ and $\{A_{i},Y_{i}\}^{N}_{i=1}$ where $M$ is the number of candidate rules and $N$ the number of observations. Specifically, for each simulated individual, we sample $\{r_{m}\}^{M}_{m=1}$ from $\{0,1\}$ with probability $\mathbf{P}_{m}\sim\operatorname{Unif}(0.1,\ 0.9)$. Then $\tau_{m}$, the subgroup CATE corresponding to a rule, is set to a fixed effect size quantity $k$ for all the true rules, and sampled from $\mathcal{N}(0,1)$ for all fake rules. To fully demonstrate the performance of D-learning under different data configurations, $k$ is set to 5,10,15, 20,25, and 30, $M$ to 100,300, and 500, and $N$ to 2500, 5000, and 10000. Accordingly, we specify the first 10, 30, and 50 rules to be the true rules; hence, $M^{*}$ varies from 10, 30, and 50. For simplicity, we pre-define the coefficient in the true linear combination, i.e., $\beta_{m}=1/M^{*}$ for all the true rules and $\beta_{m}=0$ for the fake rules. Then we can define the true ITEs based on 2 and sample $\{A_{i},Y_{i}\}^{N}_{i=1}$ in the same way in study one.
Figure 6: Number of non-zero rules selected by D-learning. Each subplot shows the result under different combinations of $\{k,M\}$ and is grouped by the number of observations NN. The gray, pink, and blue pillars show the numbers for $M^{*}=$ 10, 30, and 50, respectively.
Figure 7: Number of true rules identified by D-learning. Each subplot shows the numbers under different combinations of $\{k,M\}$ and is grouped by the number of observations NN. The gray, pink, and green pillars show the results for $M^{*}=$ 10, 30, and 50, respectively.
Together, we have $3\ (N)\times 6\ (k)\times 3\ (M^{*})\times 3\ (M)=162$ datasets in total. We apply the D-learning method on each dataset, and then compare the number of non-zero rules selected under different $M^{*}$ for 54 different combinations of $\{N,k,M\}$ to see if the aforementioned trend exists.
4.2.2 Simulation results
Figure 6 and 7 show the number of non-zero rules and true rules identified by D-learning, respectively. Clearly, across all settings, the blue pillars in 6 are higher than the pink ones and the pink than the gray, suggesting that the number of non-zero rules selected has an increasing trend with the number of true rules. Additionally, for most datasets, the number of non-zero rules identified by D-learning consistently approximates and slightly exceeds the number of true rules. As evidenced in 7, the method reliably captures most true rules while retaining a minimal set of redundant ones.
4.3 Computational complexity and efficiency of CRL
Our CRL is a multi-stage framework based on rules, and its computational cost mainly lies in the discovery, selection, and analysis of rules. Since the rule analysis step is optional and highly affected by the choice of human experts, our analysis focuses on the first two steps. Besides, we do not consider multiple runs of CRL and data splitting here in our analysis, i.e., we consider the case that all observations are used for training in both steps. In the rule generation step, suppose we use $N$ observations with $p$ covariates to train a causal forest with $T$ trees, mtry $\sqrt{p}$ and the min.node.size =s=s respectively, then we have a cost of $\mathcal{O}(T\cdot N\cdot\log(N/s)\cdot\sqrt{p})$. Subsequently, extracting rules from all the $M$ leaf nodes ($M\approx N/s$) yields a cost of $\mathcal{O}(T\cdot M\cdot\log M)$ and estimating their subgroup CATEs $\mathcal{O}(T\cdot N)$. In the rule selection step, D-learning selects a sparse subset of rules from the candidate set of size $M$, which typically yields a cost of $\mathcal{O}(M\cdot N\cdot n_{1})$ where $n_{1}\ll M$ is the number of non-zero rules selected. Together, we have a total computational cost of $\mathcal{O}(T\cdot N\cdot\log(N/s)\cdot\sqrt{p}+T\cdot N^{2}\cdot n_{1}/s)$. The $N^{2}$ term may enhance the complexity of computation, but in practice, we can mitigate this with a larger $s$, which at the same time yields shorter and thus more interpretable rules.
To better show the computational efficiency of CRL in real practice, we test CRL’s actual CPU runtimes with the data we generated in study one in section 4.1. In detail, we test three data settings: $\{N=2500,num.grp=1,k=10\}$, $\{N=5000,num.grp=3,k=20\}$ and $\{N=10000,num.grp=5,k=30\}$. In addition to the original setting of 9 covariates, we generate extra covariates with $p=18,81$ where the newly added covariates are all drawn from $\mathcal{N}(0,2)$. Therefore, we have 9 datasets in total with $N$ varying from 2500 to 10,000 and $p$ from 9 to 81. For each data set, we run CRL 100 times and report the running time quantiles. The parameters involved in CRL are fixed to their best values given in Appendix D. All computations are performed on an Apple M2, 10-core CPU, 16 GB RAM platform using R version 4.4.1.
Table 1: Running time quantiles of CRL (in seconds) based on 100 runs. NN denotes the number of observations and $p$ the number of covariates.
NN $p$ Min Q1 Mean Median Q3 Max
2500 9 0.34 0.40 0.42 0.42 0.45 0.46
18 0.37 0.43 0.45 0.45 0.48 0.49
81 0.96 0.99 1.03 1.03 1.06 1.15
5000 9 1.74 1.77 1.80 1.80 1.83 1.91
18 2.25 2.31 2.40 2.37 2.46 2.66
81 3.19 3.79 3.90 3.95 4.15 4.35
10000 9 4.35 4.41 4.56 4.47 4.73 4.93
18 4.99 5.37 5.54 5.49 5.73 6.53
81 5.29 5.42 5.75 5.57 5.80 6.78
As shown in Table 1, generally, the results are consistent with our theoretical analysis. Specifically, running time increases more than linearly with sample size $N$, as when $N$ grows from 2500 to 10,000, the increase in time is over four times ($4.35\approx 0.34\times 12$). It also increases with the number of covariates $p$. For example, when $p$ grows from 9 to 81, the increase in time is nearly about three times ($0.96\approx 0.34\times 3$). Despite the $N^{2}$ term in the theoretical cost, the actual running times remain practical (under 7 seconds even for the largest dataset), confirming that CRL is computationally feasible for the data scales commonly encountered in clinical research.